### v4

# Interpolating neural network: A lightweight yet precise architecture for data training, equation solving, and parameter calibration

Abstract
Artificial intelligence (AI) has revolutionized software development, shifting from task-specific codes (Software 1.0) to neural network-based approaches (Software 2.0). However, applying this transition in engineering software presents challenges, including low surrogate model accuracy, the curse of dimensionality in inverse design, and rising complexity in physical simulations. We introduce an interpolating neural network (INN), grounded in interpolation theory and tensor decomposition, to realize Engineering Software 2.0 by advancing data training, partial differential equation solving, and parameter calibration. INN offers orders of magnitude fewer trainable/solvable parameters for comparable model accuracy than traditional multi-layer perceptron (MLP) or physics-informed neural networks (PINN). Demonstrated in metal additive manufacturing, INN rapidly constructs an accurate surrogate model of Laser Powder Bed Fusion (L-PBF) heat transfer simulation, achieving sub-10-micrometer resolution for a 10 mm path in under 15 minutes on a single GPU. This makes a transformative step forward across all domains essential to engineering software.
1 Introduction
The evolution of software programming methodologies has transitioned from reliance on rigidly hard-coded instructions, governed by explicitly defined human rules, to the adoption of neural network-based algorithms. The transition is coined as "from Software 1.0 to Software 2.0" [1]. The shift towards Software 2.0 partially resolved the issue of labor-intensive programming in Software 1.0 and has significantly advanced the domain of large language models (LLMs) and other foundational models [2, 3]. However, the application of these technologies in the fields of engineering and science presents unique challenges. One such challenge is the relatively low generalization accuracy of machine learning (ML)-based solvers compared to traditional numerical methods [4, 5, 6, 7, 8]. After in-depth research was conducted on those methods, an increasing number of researchers began to question their effectiveness [8, 9]. Another challenge is the curse of dimensionality that arises when solving inverse design problems. As the design parameter space grows, computation requirements easily become prohibitive. Furthermore, the complexity of physical simulations continues to increase. For instance, additive manufacturing (AM) or integrated circuit (IC) simulations necessitate extremely fine resolution with high accuracy and multi-physics/multi-scale analysis within a reasonable simulation time [10, 11, 12].
In this article, we propose the concept of Engineering Software 2.0, which aims to advance the current generation of computational and data-driven engineering software into a new paradigm.
“Engineering Software 2.0 is an end-to-end software system that unifies training, solving, and calibrating in science/engineering problems."
We introduce interpolating neural networks (INNs) to achieve the new paradigm of Engineering Software 2.0. The key ideas behind INNs are: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message passing operation as a form of interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes for a given loss function. Figure 1 shows a pictorial description of the approach. When the input domain is discretized with a regular mesh (see the special case in Figure 1), an INN leverages tensor decomposition (TD) [13, 14, 15, 16] to convert the growth of the computational cost of a high-dimensional problem from exponential to linear. INN facilitates three major tasks in computational science and engineering: 1) trains on both experimental and simulation data with very high accuracy and minimal memory footprint, 2) solves high-dimensional parameterized partial differential equations (PDEs), and 3) calibrates uncertain parameters rapidly and accurately.
[…]
Figure 1: Overview of Engineering Software 2.0 enabled with Interpolating Neural Network (INN). The INN box illustrates the graph representation of the input domain discretized with an arbitrary Euclidean mesh (left) and a regular mesh as a special case (right). Node and edge features are given from the discretized mesh. After $Q$-step message passing, each node $j$ will store an interpolation function $N^{[Q](j)}(\bm{x})$. Finally, the readout operation sums the product of the interpolation functions and nodal values $\bm{u}^{(j)}$. The superscripts with square brackets [] and parentheses () denote the message passing step and graph node index, respectively. The interpolation operator $\mathcal{J}$ denotes that $\mathcal{J}\bm{u}(\bm{x})$ is a function that interpolates discrete values of $\bm{u}^{(j)}$. The INN trainer employs data-driven loss functions (e.g., mean squared error loss for regression) while the INN solver adopts a residual loss of a partial differential equation (PDE). A trained/solved INN model can then be employed as a forward model of a calibrator to solve an inverse problem.
Figure 2: Benchmark comparison of the single-scale finite difference method (FDM) solver [18], the variational multiscale FEM solver [11], and the INN solver. The data points for the first two methods with dashed marker edges are estimated, while that of INN with a solid marker edge is computed. A 3D-space, time, and 4D-parameter heat transfer equation is solved to model a 10 mm single-track laser powder bed fusion (L-PBF) metal AM. Detailed problem definition and explanation can be found in Section 3. All solvers are graphics processing units (GPU) optimized with the JAX library [19].
2 INN: Unification of Training/Learning, Solving, and Calibrating
Interpolating Neural Network (INN) generalizes numerical methods such as the finite element method (FEM) [20, 21, 22] as a subset of deep neural network (DNN) via graph network operations. Although different groups of researchers have developed the two (FEM and DNN), they share one thing in common; the goal is to find a function. The former finds a solution field of a differential equation while the latter finds an input-output relationship. INN is built upon this similarity and forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibrating, as illustrated in Figure 1.
Consider a regression problem that relates $I$ inputs and $L$ outputs. The first step of an INN is to discretize the input domain in the $I$-dimensional Euclidean space into a mesh, which can be as general as an unstructured irregular mesh or as specific as a structured regular mesh. In any case, a mesh in the Euclidean space can be readily represented as a graph – the most general form of a discretized input – as illustrated in the INN box of Figure 1. The conversion from mesh to graph has been widely accepted in the literature [23, 24, 25, 26]. Suppose there are $J$ nodes (or vertices) and $E$ edges in the graph, where the input domain is discretized with $C$ non-overlapping segments. Each segment occupies a subspace of the $I$-dimensional input domain. For instance in Figure 1, the Euclidean mesh consists of segments of hexahedrons and triangular pillars.
Each node (e.g., node $j$) has two features: nodal coordinates $(\bm{x}^{(j)}\in\mathbb{R}^{I})$ and values $(\bm{u}^{(j)}\in\mathbb{R}^{L})$. Note that the superscript with parentheses refers to the graph node index. An edge that links node $j$ and $k$ has a feature $\bm{e}^{(j,k)}$ that stores indices of segments connected to the edge. For example, when $I=3,L=2$, node $10$ in Figure 1 stores $\bm{x}^{(10)}=(x^{(10)}_{1},x^{(10)}_{2},x^{(10)}_{3})$ and $\bm{u}^{(10)}=(u^{(10)}_{1},u^{(10)}_{2})$. The edge connecting nodes $30$ and $12$ stores $\bm{e}^{(30,12)}=\{5,7\}$.
A typical message passing in a graph neural network (GNN) returns a hidden state for each graph node which is mostly a tensor (including matrix and vector) [27]. In contrast, the INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in computational mechanics) for each node $(j=1,\dots,J)$ as a hidden state. A general $Q$-step message passing of an INN, $\mathcal{M}^{[q]}(*)$, can be expressed as:
$N^{[q](j)}(\bm{x})=\mathcal{M}^{[q]}\left(\bm{x}^{(j)},\{\bm{x}^{(k)},\bm{e}^{(j,k)}|k\in\mathcal{N}^{s}(j)\},N^{[q-1](j)}(\bm{x})\right),\quad q=1,\cdots,Q,\quad N^{[0](j)}(\bm{x})=0,$ (1)
where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Supplementary Information (SI) Section 1.1 for visual illustrations). It is worth noting that the interpolation functions satisfy the Kronecker delta property, i.e., $N^{[q](j)}(\bm{x}^{(k)})=\delta_{jk}$. The operation $\mathcal{M}^{[q]}(*)$ constructs an interpolation function $N^{[q](j)}(\bm{x})$ for a graph node $\bm{x}^{(j)}$ using neighboring nodal coordinates and edge information: ${\bm{x}^{(k)},\bm{e}^{(j,k)}}$, and the interpolation function of the previous message passing: $N^{[q-1](j)}(\bm{x})$. This can be any interpolation method such as Lagrange polynomials [28, 20], meshfree [29, 30, 31], and splines [32, 33]. The $Q$-step message passing implies that it is also possible to progressively construct nonlinear or more problem-specific interpolation functions, theoretically based on the generalized finite element method [34, 35] and convolution hierarchical deep learning neural networks [7, 15]. Depending on the choice of interpolation technique, other hyper-parameters can be involved in $\mathcal{M}^{[q]}(*)$ such as the activation (or basis) function and dilation parameter [7, 15]. See SI Section 1 for various choices of the message passing operation.
During the forward propagation, an input variable $\bm{x}\in\mathbb{R}^{I}$ enters each graph node’s interpolation function $N^{[Q](j)}(\bm{x})$, followed by a graph-level readout operation:
$\mathcal{J}\bm{u}(\bm{x})=\sum^{J}_{j=1}{N^{(j)}(\bm{x})\bm{u}^{(j)}},\quad\bm{x}\in\mathbb{R}^{I},\bm{u}\in\mathbb{R}^{L},$ (2)
where the superscript $[Q]$ is dropped for brevity (i.e., $N^{[Q](j)}(\bm{x})=N^{(j)}(\bm{x})$). The interpolation operator $\mathcal{J}$ will designate an interpolated field output throughout this paper. The readout operation can be written as a tensor contraction (or matrix multiplication):
$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}=\bm{U}\bm{\mathcal{X}(x)},$ (3)
where $\bm{u}^{(j)}\in\mathbb{R}^{L},\bm{U}\in\mathbb{R}^{L\times J},\bm{\mathcal{X}(x)}\in\mathbb{R}^{J}$. Here, $\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}$ denotes the tensor contraction operation between the second dimension of tensor $\bm{U}$ and the vector (or the first dimension of tensor) $\bm{\mathcal{X}(x)}$, while $\bm{U}\bm{\mathcal{X}(x)}$ is a matrix-vector multiplication. The matrix $\bm{U}$ is a stack of nodal values $\bm{u}^{(j)}$, while $\bm{\mathcal{X}(x)}$ is a vectorized function of $\bm{x}$ that is parameterized with nodal coordinates $\bm{x}^{(j)}$ in the message passing operation.
The graph node features (i.e., coordinate ($\bm{x}^{(j)}$) and value ($\bm{u}^{(j)}$)) are trainable parameters of the INN. If the nodal coordinates ($\bm{x}^{(j)}$) are fixed, one can find nodal values ($\bm{u}^{(j)}$) without changing the discretization of the input domain. If the nodal coordinates are also updated, the optimization will adjust the domain discretization similar to r-adaptivity in FEM [21, 22]. Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, or calibrating.
Special case 1: regular mesh with Tucker decomposition
When the input domain is discretized with a regular mesh, we can significantly reduce the trainable parameters (or degrees of freedom, DoFs) by leveraging Tensor Decomposition (TD) [13]. One of the widely used TD methods is Tucker decomposition [36, 37]. It approximates a high-order tensor with $J_{i}$ nodes in $i$-th dimension as a tensor contraction between dimension-wise matrices and a core tensor $\bm{\mathcal{G}}$, which has the same order of the original tensor but with smaller nodes $M_{i}$ ($<J_{i}$). The $M_{i}$ is often called a "mode" to distinguish it from the original node $J_{i}$.
Consider a three-input ($I=3$) and one output ($L=1$) system and assume the input domain is discretized with $J=J_{1}\times J_{2}\times J_{3}$ nodes, as shown in the special case box of Figure 1. To facilitate tensor notation, the nodal values will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index, and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
The interpolated field $\mathcal{J}u(\bm{x})\in\mathbb{R}^{L=1}$ can be represented as a Tucker product:
$\mathcal{J}u(\bm{x})=\llbracket\bm{\mathcal{G}};(\mathcal{J}\prescript{}{1}{\bm{u}})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}})^{T}\rrbracket=\bm{\mathcal{G}}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}})^{T},$ (4)
where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a trainable full matrix typically smaller than the original tensor that compresses data. The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is one-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{u}^{(1)}&\prescript{1}{i}{u}^{(2)}&\cdots&\prescript{1}{i}{u}^{(J_{i})}\\ \prescript{2}{i}{u}^{(1)}&\prescript{2}{i}{u}^{(2)}&\cdots&\prescript{2}{i}{u}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{u}^{(1)}&\prescript{M_{i}}{i}{u}^{(2)}&\cdots&\prescript{M_{i}}{i}{u}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\bm{\mathcal{X}(x)}}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\bm{\mathcal{X}(x)}},$ (5)
where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}}$. As illustrated in Figure 1’s special case box, the message passing (blue arrow) only happens in the axial directions, yielding 1D interpolation functions: $N^{(J_{i})}(x_{i})$.
When there are more than one outputs $(L>1)$, the interpolated field becomes a vector of $L$ elements:
$\mathcal{J}\bm{u}(\bm{x})$ $=\left[\mathcal{J}u_{1}(\bm{x}),\mathcal{J}u_{2}(\bm{x}),\cdots\mathcal{J}u_{L}(\bm{x})\right],\textit{where}$ (6)
$\mathcal{J}u_{l}(\bm{x})$ $=\llbracket\bm{\mathcal{G}}_{l};(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T},(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T}\rrbracket=\bm{\mathcal{G}}_{l}\times^{2}_{1}(\mathcal{J}\prescript{}{1}{\bm{u}}_{l})^{T}\times^{2}_{2}(\mathcal{J}\prescript{}{2}{\bm{u}}_{l})^{T}\times^{2}_{3}(\mathcal{J}\prescript{}{3}{\bm{u}}_{l})^{T},$
where $\mathcal{J}\prescript{}{i}{\bm{u}}_{l}\in\mathbb{R}^{M_{i}\times 1}$. The trainable parameters are the core tensors $\mathcal{G}_{l}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}},l=1,\dots,L$, and the nodal values $\prescript{}{i}{\bm{U}}_{l}\in\mathbb{R}^{M_{i}\times J_{i}},l=1,\dots,L$, yielding a total count of $L\left(\prod^{I}_{i}{M_{i}}+\sum^{I}_{i}{M_{i}J_{i}}\right)$ that scales linearly with the nodal discretization $J_{i}$.
Special case 2: regular mesh with CP decomposition
Tucker decomposition can be further simplified to CANDECOMP/PARAFAC (CP) decomposition [38, 39, 40] by setting the core tensor $\bm{\mathcal{G}}$ as an order-$I$ super diagonal tensor: $\bm{\mathcal{G}}\in\mathbb{R}^{M^{I}}$, $M=M_{1}=\cdots=M_{I}$, all zero entries except the diagonal elements. If we further set the diagonal elements of $\bm{\mathcal{G}}$ to be 1, the Tucker decomposition in Eq. 6 becomes:
$\mathcal{J}\bm{u}(\bm{x})=\sum^{M}_{m=1}{[\mathcal{J}\prescript{m}{1}{\bm{u}}(x_{1})\odot\dots\odot\mathcal{J}\prescript{m}{I}{\bm{u}}(x_{I})]},$ (7)
where $\mathcal{J}\prescript{m}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{L}$ and $\odot$ represents element-wise multiplication. In CP decomposition, the core tensor is no longer trainable, thus the total trainable parameter becomes $ML\sum^{I}_{i}{J_{i}}$.
It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.5 is one-dimensional. In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs). As given in Table 1, the number of trainable parameters of the full interpolation scales exponentially with the discretization $J_{i}$, whereas the Tucker and the CP decomposition scale linearly with $J_{i}$. Considering the trainable parameters of multi-layer perception (MLP) scales quadratically with the number of hidden neurons, INNs with tensor decomposition (TD) may dramatically reduce the model complexity and computing requirements.
Table 1: Number of trainable parameters (or degrees of freedom, DoFs) of full interpolation, Tucker decomposition, and CP decomposition for an $I$-input single output $(L=1)$ relationship. The input domain is discretized with a regular grid of $J_{1}\times J_{2}\times\cdots\times J_{I}$ discretization. We assume that the nodal coordinates $(\bm{x}^{(j)})$ are fixed.
Full interpolation Tucker decomposition CP decomposition
Trainable parameters $\prod^{I}_{i}{J_{i}}$ $\prod^{I}_{i}{M_{i}}+\sum^{I}_{i}{M_{i}J_{i}}$ $M\sum^{I}_{i}{J_{i}}$
INN trainer
A general INN forward propagation is provided in Eq.3 as a tensor contraction. The INN trainer optimizes nodal values $\bm{U}=\{\bm{u}^{(j)}\}_{j=1,\cdots,J}$ (and if needed, nodal coordinates ($\bm{X}=\{\bm{x}^{(j)}\}_{j=1,\cdots,J}$)) under a given loss function and training data. Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$; and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$. The superscript * denotes data. A mean squared error (MSE) loss function for this regression problem is defined as:
$loss(\bm{U},\bm{X})=\frac{1}{K}\sum_{k}(\mathcal{J}\bm{u}(\bm{x}^{*}_{k})-\bm{u}^{*}_{k})^{2}.$ (8)
Finally, an optimization is formulated as:
$\underset{\bm{U},\bm{X}}{\text{minimize}}\ loss(\bm{U},\bm{X}).$ (9)
Section 3.1 exhibits the application of INN trainer to metal additive manufacturing while other benchmarks are elucidated in SI Section 2.
INN solver
An INN solver generalizes classical numerical methods such as FEM and meshfree, as well as the model order reduction methods such as proper generalized decomposition (PGD) [41] and tensor decomposition (TD) [14, 42]. Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ (S-T-P) PDE, whose computational cost is prohibitively high for most numerical methods and machine learning approaches [41].
Consider a parameterized space-time PDE:
$\mathcal{L}\bm{u}(\bm{x},t,\bm{\theta})=0,$ (10)
where $\mathcal{L}$ is the general partial differential operator (can be linear or nonlinear). The INN solver uses the same neural network structure as the trainer, but the loss function varies with the equations to be solved. It is formulated as the weighted summation of the PDE residual [43, 44, 45]. An INN solution field in the S-T-P domain can be written as:
$\begin{split}\mathcal{J}{\bm{u}}(\bm{x},t,\bm{\theta};\bm{U},\bm{X})&=\sum_{m=1}^{M}[\mathcal{J}\prescript{m}{x_{1}}{\bm{u}}(x_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{x_{d}}{\bm{u}}(x_{d})]\odot[\mathcal{J}\prescript{m}{t}{\bm{u}}(x_{t})]\\ &\odot[\mathcal{J}\prescript{m}{\theta_{1}}{\bm{u}}(\theta_{1})]\odot\cdots\odot[\mathcal{J}\prescript{m}{\theta_{k}}{\bm{u}}(\theta_{k})],\end{split}$ (11)
where $d$ and $k$ are the spatial dimension and the number of parameters, respectively. Similar to the trainer, the goal is to find nodal values $\bm{U}$ (and nodal coordinates $\bm{X}$, if one wants to adapt the mesh). As a result, the INN obtains the S-T-P solution by minimizing the loss function:
[…]
where $\delta\bm{u}$ is the weight function and can be defined using Galerkin, Petrov-Galerkin, collocation or quadratic formulation. Due to the Kronecker delta property of INN interpolation functions, Dirichlet boundary conditions and initial conditions can be strongly imposed. As a result, only the weighted sum residual is considered in Eq. 11 as a loss function. This distinguishes INNs from most data-driven PDE solvers that weakly impose those conditions [46]. See the SI Section 4 for detailed derivations of the loss function and solution scheme. INN solvers applied to the metal additive manufacturing and the linear elastic solid mechanics problem are introduced in Section 3.3 and SI Section 4, respectively.
INN calibrator
[…]
$\bm{x}^{*}=\underset{\bm{x}}{\text{argmin}}\frac{1}{K}\sum_{k}(\mathcal{J}\bm{u}(\bm{x})-\ \bm{u}^{*}_{k})^{2},$ (13)
where $\mathcal{J}\bm{u}(\bm{x})$ is the trained or solved forward model, $\bm{u}^{*}_{k}$ is the $k$-th measured observation and $\bm{x}^{*}$ is the calibrated input variable. The INN calibrator applied to the heat source calibration task in metal additive manufacturing can be found in Section 3.3 while other benchmarks can be found in SI Section 3.
3 Scientific and Engineering applications: Metal Additive Manufacturing
Figure 3: A summary of engineering applications of INNs to additive manufacturing (AM) problems. First, a data-driven real-time online monitoring and feedback control tool is formulated with an INN that uses only 18% training parameters of MLP and is 18-31 times faster than MLP. The calibration problem develops a reduced-order model of laser powder bed fusion (L-PBF) AM to calibrate the heat source parameters from experimental data. Finally, the INN solver solves a space-time-parameter heat transfer equation, resulting in a significant storage reduction and faster simulations.
This section illustrates how Engineering Software 2.0 enabled with INNs can be applied to advance various facets of metal additive manufacturing (AM). Note that the INN with CP decomposition will be referred to as INN in this section for brevity. Laser powder bed fusion (L-PBF) is a type of metal additive manufacturing in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [47]. The added flexibility of L-PBF comes with a huge design space and spatially varying microstructure after manufacturing. Therefore, considerable research is devoted to employing in situ monitoring data for real-time control of L-PBF processes to minimize the variability and optimize the target properties [48]. Computational modeling of L-PBF spans the simulation of the manufacturing process with a high dimensional design space (solving), the identification of model parameters from sparse experimental data (calibrating), and the online monitoring and control of the L-PBF (training).
The governing equation for modeling the manufacturing process is given by a heat conduction equation with a moving laser heat source $S_{h}$:
$\frac{\partial T(\mathbf{x},t)}{\partial t}-k\Delta T(\mathbf{x},t)=S_{h}(\mathbf{x},t,P,\eta,d)$ (14)
where the equation involves spatial variables $\mathbf{x}$, temporal variable $t$, and four variable parameters: thermal conductivity $k$, laser power $P$, absorptivity $\eta$, and laser depth $d$. For the simplest case of single track laser pass, the initial condition is $T(\mathbf{x},0)=T_{0}(\mathbf{x})$, and the boundary conditions (BCs) are $T=\tilde{T}$ in $\Gamma^{D}$ for Dirichlet BC and $\frac{\partial T}{\partial\mathbf{x}}\cdot\mathbf{n}=\bar{q}$ in $\Gamma^{N}$ for Neumann BC, where the spatial domain boundary $\Gamma=\Gamma^{D}\cup\Gamma^{N}$.
3.1 Real-time Online Control of Laser Powder Bed Fusion Additive Manufacturing
The first application is fully data-driven, where the data are generated from numerical methods and the INN trainer is used to develop an online control system for L-PBF AM. The goal is to maintain a homogeneous temperature of the melt pool (i.e., a molten region near the laser spot) across each layer using real-time sensor data, under the hypothesis that a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties. The laser power at each point of a certain layer needs to be controlled (increased/decreased by certain amount compared to previous layer) to achieve this homogeneity in melt pool temperature. The fundamental challenge of implementing a model predictive control system for such an application is the computational resources required for forward prediction (involving finite element or computational fluid dynamics models) and inverse prediction, as illustrated in Figure 3. The INN trainer aims to provide a reliable and memory-efficient surrogate model that can replace mechanistic models in predictive control systems. In this application, the database is generated by fusing experiments and computational methods (see SI Section 2.6 for details) for aluminum alloy. The online control loop uses the trained INN models to inform the manufacturing machine. There are two models: forward model and inverse model.
The forward model uses the mechanistic features and the thermal emission plank (TEP, a representative of the melt pool temperature) to predict the TEP of the next layer of the build when no correction is applied. The mechanistic features are precomputed based on the tool path and geometry of the build to generalize a data-driven model for unseen tool path and geometric features [49]). The target TEP for the next layer is the average of the forward model’s outputs (i.e., predicted TEPs for the next layer) measured at the current layer. The inverse model takes this target TEP as the input along with the previous layer’s mechanistic features to predict the required laser power for the next layer. Finally, the control software calculates how much change in laser power is required at each point, and sends the information to the machine. The corrected laser power is then used to print the next layer.
The forward model is trained on experimentally observed data, while the inverse model is trained on computational data coming from a finite-difference solver [18]. The INN model is applied for both forward and inverse prediction and is compared with the MLP. Figure 3 shows that for the inverse model INN is at least 18 to 31 times faster to train compared to the MLP to reach the same level of training error, and it reduced the number of trainable parameters by a staggering amount of 82%. The speed-up depends on the amount of data used. The INN method is consistently more accurate compared to MLP (see, SI Section 2.6 for more details).
3.2 Calibration of Heat Source Parameters for Additive Manufacturing
In this example, an INN calibrates the heat source function $S_{h}$ in Eq.14 for experimental data. A Gaussian beam profile is modeled as a volumetric heat source [50], which is written as:
$S_{h}(\mathbf{x},t,\mathbf{p})=\frac{2P\eta}{\pi r^{2}_{b}d}e^{-\frac{2(x^{2}+y^{2})}{r^{2}_{b}}}\cdot\mathbf{1}_{\{z|z_{top}-z\leq d\}}(z)$ (15)
where $z_{top}$ is the z-coordinate of the current surface of the powder bed and $\mathbf{1}$ is the indicator function. The notation follows Eq. 14, and $r_{b}$ is the Gaussian profile’s standard deviation that controls the beam’s width. Following reference [10], the heat source parameters $d,\eta,r_{b}$ are controlled by the calibration parameters $p_{1}$, $p_{2}$, $p_{3}$ such that $d=p_{1}\frac{P}{V}RHF^{2}$, $\eta=p_{2}\frac{P}{V}RHF^{2}$, and $r_{b}=p_{3}\frac{P}{V}RHF^{2}$. Here, $V$ is the laser scan speed and $RHF$ is a predetermined residual heat factor based on the laser toolpath [51].
This calibration problem’s major challenge is achieving a highly accurate surrogate model given a sparse set of training data. INNs outperform MLPs in solving these data-scarce problems because they are rooted in the interpolation theory. As a demonstration, an INN with 1760 trainable parameters is trained on 28 sets of meltpool temperature data generated by the finite difference method (FDM) [18] using 4 different power-scanning speed set-ups. Later, this trained INN is used as a surrogate to optimize the three parameters ($p_{1},p_{2},p_{3}$) given the experimental data. To show the superiority of INNs over traditional neural networks in such data-scarce applications, a 3-layer MLP with 2689 trainable parameters (roughly the same order of magnitude as the INN) is trained using the same data fed to the INN. In practice, we observed that INN converges well within 100 epoch, while MLP requires over 1000 epoch to converge to a test MSE 10 times larger than INN given all other setup the same. Details on this training and calibration process are provided in SI Section 3.1.
The INN- and MLP-based calibrations are repeated 50 times to provide a statistical comparison (see Figure 3). Even with only 28 sets of simulation data, the calibrated parameters using the INN produce a mean melt pool temperature within a 7.9% difference of the experimental data. In contrast, the parameters calibrated from the MLP surrogate produce a difference of 44. 1%, which is unsatisfactory. This is expected since MLP is a general model requiring a large amount of data to achieve a reliable representation of the problem. On the other hand, the INN achieves a much lower difference, showing that the model can learn underlying physics even with a scarce dataset.
3.3 Solving High Dimensional Space-Time-Parameter Heat Equation
[…]
The surrogate model $f$ can also be obtained using the classical data-driven approach. In the offline stage, training data can be generated by repeatedly running numerical solvers (such as finite difference method (FDM) and variational multiscale method (VMS)) in the Space-Time domain with different sets of parameters $\mathbf{{p}}$. Then a surrogate model is trained with the data. The standard data-driven approach suffers from the curse of dimensionality when the parametric space is high-dimensional. As a result, the data-driven approaches face three challenges at the offline stage: expensive computation for repetitive data generation, huge memory cost for running full-scale simulations, and excessive disk storage of offline training data. We estimate the time and storage required for the data-driven approach with FDM and VMS, as shown in Table 2 and illustrated in Figure 2.
Unlike the data-driven approaches, INN solvers treat the parameters $\mathbf{p}$ as additional parametric inputs. As a result, INNs can directly obtain a parametric surrogate model $f=u(\mathbf{x},t,\mathbf{p})$ directly from the governing equation without going through the cumbersome offline data generation and surrogate model training process. This is particularly beneficial to high-dimensional parametric surrogate models. Moreover, our results prove the INN solution (or the S-T-P surrogate model) achieves an exceptionally high accuracy of $R^{2}$ = 0.9969, considering that most data-driven approaches for S-T-P problems trained on numerical simulation data suffer from low model accuracy, typically below $R^{2}=0.9$ [52]. It is worth mentioning that the physics-informed neural networks (PINNs) can handle the S-T-P problem similar to the INN solver. However, the number of collocation points scales exponentially with the number of inputs, and our preliminary study revealed that PINN fails to converge for the same 8-dimensional S-T-P problem after consuming considerable computational resources (see SI Section 4.4). The performance comparison of INN versus data-driven methods is shown in Table. 2.
Table 2: Performance comparison of different surrogate models. The INN’s discretization in the 4D parametric space is $100^{4}$, so the corresponding sampling points required for the data-driven methods should be $10^{8}$. The total simulation time and storage requirement of data-driven approaches are estimated from a single run. Detailed explanations for this benchmark can be found in SI Sections 4.3 and 4.4.
Method
Resolution
($\mu$m)
Number of
offline simulations
Total offline
simulation time
Storage
Data-driven (FDM) 25 $1\times 10^{8}$ 2,600 years $7.82\times 10^{8}$
Data-driven (VMS) 12.5 $1\times 10^{8}$ 9 years $8.79\times 10^{6}$
INN 8 $1$ 15 minutes $7.37\times 10^{-2}$
4 Conclusions
This article demonstrates that the interpolating neural network (INN) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods and machine learning models. The keys to INN’s success are 1) it interpolates the graph nodes’ values with well-established interpolation theories (not interpolating training data with a black-box network that often causes overfitting issues) and 2) it leverages the tensor decomposition (TD) to resolve curse of dimensionality. Due to the significantly reduced trainable parameters without losing approximation capabilities, INN can train/solve/calibrate faster with less memory requirement, and achieve extremely high model accuracy and fast inference time.
As this is the original article on INNs, we have myriads of research questions that may gather huge research interests across various areas including but not limited to scientific machine learning, applied mathematics, computer vision, and data science. To mention a few:
  • Depending on the problem, INN’s superior training efficiency may cause overfitting issues. The ensemble training (i.e., training on random subsets of the training data) might be a way to enhance the generality of the model for image problems or noisy data.
  • Multi-resolution aspect (i.e., a varying number of segments across TD modes) of INNs needs to be investigated. See SI 1.4 for further discussions.
  • Similar to the INN solvers, the convergence of INN trainers can be studied mathematically.
  • Since INNs can be used as both a solver for physics-based problems and a function approximator for data-driven problems, it is of interest to combine the two behaviors into one single model to solve large-scale problems involving both complex physics and scarce data.
  • The superior performance of the INN solver may facilitate multiscale analysis in a vast parametric space. One can integrate the S-T-P INN solver with a concurrent multiscale analysis framework such as self-consistent clustering analysis [53, 54] and open a new direction towards parameterized multiscale analysis.
  • The current INN code has been optimized to run on a single GPU. Since the INN forward pass over multiple modes can be parallelized, multi-GPU programming will further speed up the code.
  • Although Tucker decomposition has been introduced as a special case 1 in Section 2, the numerical experiments conducted in this article mostly focus on the CP decomposition. Theoretically, Tucker decomposition has a larger approximation space than CP decomposition. Further studies on the numerical aspects of the Tucker decomposition are needed.
We expect INNs will open a next-generation Engineering Software 2.0 that revolutionizes almost all fields of computational science and engineering.
Acknowledgments
We express our sincere gratitude to Dr. Gino Domel, Mr. Joseph P. Leonor, Mr. Stefan Knapik, and Mr. Vispi Karkaria from Northwestern University for their invaluable comments and support throughout this work. We also acknowledge the following grants: DEVCOM Army Research Laboratory under Cooperative Agreement Numbers W911NF-20-2-0292 and W911NF-21-2-02199 to J. Cao and the startup funds provided by Virginia Polytechnic Institute and State University to S. Saha. C. Park acknowledges the internship opportunity with the Freudenberg Group, within the team managed by Dr. Vishwanath Hegadekatte.